OpenStudy (anonymous):

volume of solid enclosed by surface (x^2+y^2+z^2)^2=4z(x^2+y^2)

OpenStudy (anonymous):

x=0, y=(-3sqrt3)/4 ,z=3/4 x=0, y=(3sqrt3)/4 ,z=3/4